Solve the equation. Points: 6 3x +6+6-11
Answer #1

|3x + 6 | + 6 = 11

subtracting both sides by 6

| 3x + 6 | = 5

removing the absolute value bars

3x + 6 = +- 5

3x + 6 = 5

x = -1/3

3x + 6 = -5

x = -11/3

solution set is { -1/3 , -11/3 }